1.     Assume that the market demand and the costs of the duopolists are:

P=120-0.4(QA + QB)

TCA=5QA

TCB= 0.2Q2B

Also assume that firm B is the sophisticated leader, Determine:


1.     The reaction curve of A

2.     The reaction curve of B

3.     The profit function of A

4.     Stackelberg equilibrium output level for firm A

5.     Stackelberg equilibrium output level for firm B

6.     The market price.


1
Expert's answer
2021-08-30T14:36:53-0400

Given Information

"P=120-0.4(Q_A + Q_B)\\

TCA=5Q_A\\

TCB= 0.2Q^2_B"


Calculate Marginal Revenue

For Firm A

"P=120-0.4(Q_A + Q_B)"


"TR =Price\times Quantity"

"TR = (120-0.4(Q_A + Q_B)\times Q_A\\

TR = 120\times Q_A - 0.4\times Q_A^2 - 0.4Q_AQ_B"

Now derivate TR with respect to Quantity we get,

"MR = 120 - 0.8Q_A- 0.4Q_B"

 

For Firm B

"P=120-0.4(Q_A + Q_B)"

"TR =Price\times Quantity"

"TR = (120-0.4(Q_A + Q_B)\times Q_B\\

TR = 120\times Q_B - 0.4\times Q_AQ_B- 0.4Q_B^2"

Now derivate TR with respect to Quantity we get,

"MR = 120 - 0.4Q_A- 0.8Q_B"


Calculate Marginal cost for both the firms.

For Firm A

"TCA=5Q_A"

derivate TC with respect to quantity to calculate MC, We get

"MC = 5\\

TCB= 0.2Q^2_B"

derivate TC with respect to quantity to calculate MC, We get

"MC = 0.4Q_B"


1)

For best response function or for reaction function put MR = MC

For Firm A

"MR_A − MC_A\\

120 - 0.8Q_A- 0.4Q_B = 5\\

0.8Q_A+ 0.4Q_B =115\\

0.8Q_A =115 - 0.4Q_B\\

Q_A=143.75 - 0.05Q_B"

Reaction Function for Firm A is "Q_A=143.75 - 0.05Q_B"


2)

For Firm B

"MR_B − MC_B\\

120 - 0.4Q_A- 0.8Q_B = 0.4Q_B\\

1.2Q_B = 120 - 0.4Q_A\\

Q_B = 100 - 0.333Q_A"

Reaction Function for Firm B is "Q_B = 100 - 0.333Q_A"


3)

Profit Function of A

"Profit = TR_A - TC_A\\

Profit for A = 120\times Q_A - 0.4\times Q_A^2 - 0.4Q_AQ_B - 5Q_A\\

Profit for A = 115Q_A - 0.4\times Q_A^2 - 0.4Q_AQ_B"


4)

The Stackelberg leader will choose its output QA to Max its profits, s.t. reaction function of the firm B

"π_A =120\times Q_A - 0.4\times Q_A^2 - 0.4Q_A\times (100 - 0.333Q_A) - 5Q_A\\

π_A = 115Q_A - 0.4\times Q_A^2 - 0.4Q_A\times (100 - 0.333Q_A)\\ 

π_A = 115Q_A - 0.4\times Q_A^2 - 40Q_A + 0.133Q_A\\

π_A = 75.132Q_A - 0.4\times Q_A^2"

derivate function with respect to QA and put equals to 0, we get

"75.132 - 0.8\times Q_A = 0\\

75.132 = 0.8\times Q_A\\

Q_A = 93.9165"


5)

Put Value of Q = 93.9165 in Equation "Q_B = 100 - 0.333Q_A"  we get

"Q_B = 100 - 0.333Q_A\\

Q_B = 100 - 0.333\times (93.9165)\\

Q_B = 100 - 31.3023\\

Q_B = 68.697"


6)

market price

"P=120-0.4(Q_A + Q_B)"

Put value Q = 93.9165 and QB = 68.697, we get

"P=120-0.4(93.9165 + 68.697)\\

P=120-0.4(162.6141)\\

P=120 - 65.0456\\

P=54.9543"